OpenStudy (anonymous):

Solve by substitution: 8x-3y=17 8x-y=11

OpenStudy (angela210793):

subtract both eq and U'll get -2y=6--->y=-3 replace -2 in the 2nd or 1st eq to find x 8x+3=11 8x=8 x=8/8=1 solution(1:-3)

OpenStudy (anonymous):

from 2nd we have: y=8x-11 8x-3(8x-11)=17 8x-24x+33=17 -16x=-16 16x=16 x=1

OpenStudy (amistre64):

substitution is sooo labor intensive, id rather eliminate... 8x-3y=17 ; *-1 8x- y=11 -8x +3y =-17 8x - y = 11 ------------- 2y = -6 y = -3